OpenStudy (anonymous):

WILL MEDAL If cosine of x equals 1 over 2, what is sin(x) and tan(x)? PLEASE explain your steps.

Nnesha (nnesha):

\[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

 cos =adjacent over hyp so write that and then apply Pythagorean theorem to find 3rd side of right triangle .

OpenStudy (anonymous):

tan=?/1 sin=?/2

OpenStudy (anonymous):

1^2+2^2=c^2?

Nnesha (nnesha):